OpenStudy (anonymous):

Use a proof by cases to show that min(a , min(b, c» = min(min(a , b) , c) whenever a, b, and c are real numbers.

OpenStudy (anonymous):

i guess you can grind it out with cases, for example if \(a\leq b\leq c\) then \[\min(a, (\min(b,c))=\min(a, b)=a\]

OpenStudy (anonymous):

similarly if \(a\leq c\leq b\) then \(\min(a,\min(b,c))=\min(a,c)=a\) and so on

OpenStudy (anonymous):

thank you for clear answer

OpenStudy (anonymous):

oh i guess you have to say for the first one that \(\min(a, \min(b,c))=a=\min(\min(a,b),c)\)

OpenStudy (anonymous):

likewise for the second. there may be a snappier way to do this, but it did say by cases, so just grind it out

OpenStudy (anonymous):

yw